# EvolveNav: Self-Improving Embodied Reasoning for LLM-Based Vision-Language Navigation

### v2

@@ Abstract
Building Vision-Language Navigation (VLN) agents which can navigate following natural language instructions is a long-standing goal in human-robot interaction applications. Recent studies have revealed the potential of training open-source Large Language Models (LLMs) to unleash LLMs’ reasoning ability for improving navigation, and simultaneously mitigate the domain gap between LLMs’ training corpus and the VLN task. However, these approaches primarily adopt direct input-output mapping paradigms, causing the mapping learning difficult and the navigational decisions unexplainable. Chain-of-Thought (CoT) training is a promising way to improve both navigational decision accuracy and interpretability, while the complexity of the navigation task makes the perfect CoT labels unavailable and may lead to overfitting through pure CoT supervised fine-tuning. In this paper, we propose a novel sElf-improving embodied reasoning framework for boosting LLM-based vision-language Navigation, dubbed EvolveNav. Our EvolveNav consists of two stages: (1) Formalized CoT Supervised Fine-Tuning, where we train the model with formalized CoT labels to both activate the model’s navigational reasoning capabilities and increase the reasoning speed; (2) Self-Reflective Post-Training, where the model is iteratively trained with its own reasoning outputs as self-enriched CoT labels to enhance the supervision diversity. A self-reflective auxiliary task is also introduced to encourage learning correct reasoning patterns by contrasting with wrong ones. Experimental results on the popular VLN benchmarks demonstrate the superiority of EvolveNav over previous LLM-based VLN approaches. Code is available at: https://github.com/expectorlin/EvolveNav.
@@ 1 Introduction
Vision-Language Navigation (VLN) has received significant research interest within the Embodied AI community, due to its practicality and flexibility in enabling human-robot interaction in real-world robotic applications. In VLN tasks, an embodied agent needs to follow natural language instructions to navigate through complex visual environments to reach the target position. Early works improve VLN performance by designing dedicated model architectures wang2019reinforced ; ma2019self ; deng2020evolving ; qi2020Object , introducing powerful learning paradigms zhu2020vision ; li2019robust ; huang2019transferable , and developing useful data augmentation techniques tan2019learning ; fried2018speaker ; liu2021vision ; Fu2019CounterfactualVN . Subsequently, pretraining-based VLN approaches have been widely proposed to improve the cross-modal alignment ability and decision accuracy of navigation agents hong2021vln ; Chen2021HistoryAM ; Qi2021TheRT ; Chen2022ThinkGA ; Qiao2022HOPHA . Nevertheless, constrained by the limited scale of pretraining and VLN in-domain data, these approaches cannot learn navigational reasoning knowledge sufficiently, and therefore still struggle to handle various unseen navigation scenarios.
With the rapid progress of large language models (LLMs) brown2020language ; touvron2023llama ; touvron2023llama2 , emerging works have introduced LLMs to address embodied tasks considering LLMs’ rich real-world common sense and powerful reasoning ability Ahn2022DoAI ; Huang2022InnerME ; driess2023palme . Some recent works have attempted to build LLM-based VLN models in a zero-shot or trainable manner  chen2024mapgpt ; zhou2023navgpt ; long2023discuss ; hao2020towards ; zhang2024navid . The zero-shot approaches, such as NavGPT zhou2023navgpt and MapGPT chen2024mapgpt , resort to closed-source LLMs OpenAI_2023 to generate navigational reasoning and decision for different navigation timesteps. To alleviate the high cost of frequently querying closed-source LLMs for sequential decision making, the trainable methods collect in-domain data to train open-source LLMs vicuna to build the navigation agent. However, they typically map directly from navigational inputs to decisions without explicit intermediate reasoning steps, leading to decision uninterpretability and may also limit performance (see Figure 1(a)).
@@
Recent studies have revealed the effectiveness of embodied chain-of-thoughts (CoT) training in enhancing both the decision accuracy and interpretability for embodied tasks zawalski2024robotic ; zhao2025cot ; lin2024navcot . Most of these approaches employ the supervised fine-tuning paradigm for embodied CoT training. However, introducing CoT supervised fine-tuning for training VLN models is challenging. Firstly, due to the complexity and uncertainty of the navigation task, there can be multiple cues for deciding the correct navigation action, i.e., there can be multiple “correct” CoT supervisions to guide navigation for a specific timestep. This leads to a hard collection process of perfect navigation CoT labels. Secondly, pure CoT supervised fine-tuning using fixed CoT labels may cause overfitting to certain reasoning patterns and harm the generalization to diverse unseen scenarios.
@@
In this paper, we propose a sElf-improving embodied reasoning paradigm for enhancing LLM-based vision-language Navigation, called EvolveNav, to fulfill generalizable and adaptable navigational reasoning under different scenes. Our EvolveNav comprises two stages: 1) Formalized CoT Supervised Fine-Tuning and 2) Self-Reflective Post-Training. The Stage 1 training of Formalized CoT Supervised Fine-Tuning aims to first activate the model’s potential reasoning capabilities, where we ask the model to produce explicit chain-of-thought navigational reasoning by predicting the landmarks needed to locate with the corresponding direction for deciding the navigation actions. To alleviate redundant reasoning and increase the inference speed, we conduct the CoT supervised fine-tuning using formalized CoT labels, which are collected by filling the landmark and direction information into concise label templates. To mitigate the overfitting to pre-constructed CoT labels and enable self-improving reasoning for enhancing generalization, we further conduct Self-Reflective Post-Training for Stage 2 training. Concretely, we design a self-enriched CoT label enhancement strategy, where we train the model with its iteratively produced correct reasoning outputs to enrich the CoT supervision. We also introduce a self-reflective auxiliary task, where the model needs to discriminate between positive and negative navigational reasoning to learn correct reasoning patterns better. As shown in Figure 1, in contrast to CoT supervised fine-tuning using fixed CoT labels (i.e., Figure 1(b) Formalized CoT Learning and (c) Free-form CoT Learning), our EvolveNav (Figure 1(d)) can generate embodied CoT in a self-refining manner during training to mitigate the overfitting. Additionally, through training with formalized CoT labels, our EvolveNav can significantly reduce uninformative navigational reasoning to promote the reasoning speed compared with using free-form CoT labels for training (as in Figure 1 (c)).
@@
We conduct experiments on two public VLN benchmarks, R2R anderson2018vision and CVDN thomason2019vision . Experimental results show that EvolveNav substantially outperforms previous LLM-based VLN approaches on unseen scenarios, demonstrating the effectiveness of our self-improving embodied reasoning paradigm in promoting navigation decision accuracy and generalization. We carefully conduct ablation experiments to explore how to design streamlined CoTs that can provide interpretability while boosting navigation performance. Visualization also insightfully reveals the reasonability of our design for CoT labels in enhancing decision interpretability and improving navigational reasoning.
@@
  • We propose a novel self-improving embodied reasoning paradigm, EvolveNav, for enhancing the LLM-based vision-and-language navigation, which fulfills generalizable and adaptable navigational reasoning under unseen scenes.
@@
  • Experimental results show the superiority of EvolveNav over other LLM-based approaches on two standard VLN benchmarks R2R and CVDN. Our EvolveNav can improve the generalization of both navigational reasoning and decision-making, providing meaningful insights for designing further embodied reasoning paradigms.
@@ 2 Related Work
Vision-Language Navigation (VLN). Vision-Language Navigation (VLN) has attracted intensive research interest in recent years. Various VLN benchmarks have been proposed to evaluate agents’ ability for navigational reasoning and instruction following anderson2018vision ; jain2019stay ; ku2020room ; thomason2019vision ; qi2020reverie ; zhu2021soon . Previous approaches employ non-pretraining-based wang2019reinforced ; ma2019self ; deng2020evolving ; zhu2020vision ; tan2019learning ; fried2018speaker or pretraining-based paradigms hong2021vln ; Chen2021HistoryAM ; Chen2022ThinkGA ; Qiao2022HOPHA ; Guhur2021AirbertIP ; an2022bevbert ; wang2023scaling for tackling the above VLN tasks. However, these approaches cannot generalize well to diverse unseen scenarios that require rich real-world commonsense, and the navigation decisions also lack explainability. Some recent works have introduced LLMs to assist the VLN task, by either eliciting the useful navigation knowledge stored in LLMs lin2024correctable ; zhou2024navgpt ; qiao2023march or employing the LLM as the navigation backbone for action decision zhou2023navgpt ; zheng2024towards ; zhang2024navid ; lin2024navcot ; chen2024mapgpt . Our work lies in the latter. Different from previous approaches, we propose a new VLN framework in this work, where the LLM-based navigation backbone iteratively generates intermediate reasoning steps in a self-improving manner during training to guide navigational decisions. As a result, both the reasoning ability and decision interpretability of the navigation model can be significantly enhanced.
@@
LLMs as Embodied Agents. Recent research have revealed the giant potential of utilizing Large Language Models (LLMs) as embodied agents to complete the robotic navigation and manipulation tasks, benefiting from the outstanding ability of planning, reasoning, and reflection of LLMs song2023llmplanner ; Ahn2022DoAI ; Huang2022InnerME ; Yao2022ReActSR ; shahlm ; schumann-2023-velma ; wang2023voyager ; huang2023voxposer ; rajvanshi2024saynav . For example, LM-Nav shahlm introduces the LLM to parse the long navigation instruction into sequential landmarks for facilitating the navigational planning. Voxposer huang2023voxposer introduces LLMs for code writing and combines them with the Vision-Language models (VLMs) to compose 3D value maps for robotic manipulation.
@@
There are typically two branches of works where the LLMs act as embodied agents for tackling the VLN task. In the first branch, closed-source LLMs like GPT-4 openai are queried in a zero-shot manner to decide the action sequentially zhou2023navgpt ; long2023discuss ; chen2024mapgpt ; chen2024affordances . For example, NavGPT zhou2023navgpt transforms visual observations into textual formats and feed them to LLMs for generating action predictions. The second branch finetunes open-source LLMs with in-domain VLN datasets, which alleviates the LLM’s query cost as well as mitigates the gap between LLM’s training corpus and VLN tasks zheng2024towards ; lin2024navcot ; zhang2024navid ; han2024roomtour3d . For example, Navid zhang2024navid constructs a video-based navigational vision-language model and train it using navigation samples collected from continuous R2R datasets. However, most of them map navigation inputs to action decisions without the reasoning output. In contrast, we train the open-source LLM to generate self-improving embodied reasoning explicitly to improve action decision accuracy, which enhances the decision interpretability as well as mitigates overfitting to training reasoning labels.
@@
Embodied Chain-of-thoughts Training. Chain-of-thoughts (CoT) reasoning wei2022chain has been a widely utilized technique in Large Language Models (LLMs) and Vision-Language Models (VLMs). By generating the intermediate reasoning steps rather than directly predicting the answer, CoT can promote the answer accuracy for various tasks such as mathematical reasoning, commonsense reasoning, code generation, etc openai ; OpenAI_2023 ; liu2024deepseek . Inspired by this, some recent works have trained LLMs/VLMs to generate reasonable CoT for improving the action decision accuracy in embodied tasks zawalski2024robotic ; lin2024navcot ; mu2023embodiedgpt ; zhao2025cot ; liu2025spatialcot ; zhou2024navgpt . ECoT zawalski2024robotic trains vision-language-action models to generate embodied reasoning including object bounding boxes and end effector positions to encourage better adaptation for robotic manipulation tasks. NavGPT-2 zhou2024navgpt resorts to the GPT-4V model to collect free-form step-wise CoT reasoning data to improve the navigational reasoning ability of the VLN-specialized models Chen2022ThinkGA . CoT-VLA zhao2025cot incorporates explicit visual chain-of-thought reasoning into vision-language-action models by predicting future image frames before generating the action sequence. In our work, we propose a novel self-improving embodied CoT training paradigm for boosting the VLN, which effectively improves the reasoning ability and decision accuracy of the navigation model in various unseen scenarios. Moreover, we construct formalized CoT supervision, which can reduce redundant reasoning information and simultaneously promote the model’s inference speed.
@@
Figure 2: Overview of our EvolveNav. EvolveNav introduces a two-stage training framework for fulfilling self-improving embodied reasoning. In Stage 1 Formalized CoT Supervised Fine-Tuning, the navigation agent is trained using pre-constructed formalized CoT labels to generate navigational reasoning by predicting the landmark needed to locate with the corresponding direction. In Stage 2 Self-Reflective Post-Training, the agent’s own reasoning outputs are introduced as the self-enriched CoT labels to enhance the supervision diversity. A self-reflective auxiliary task is also designed to guide the navigation agent to discriminate between correct and wrong reasoning outputs.
@@ 3 Method
In this section, we first introduce the problem definition of the VLN task (Sec. 3.1). Then, we present the model architecture of the LLM-based navigation agent in our EvolveNav (Sec. 3.2). Finally, we describe our proposed self-improving embodied reasoning framework in detail (Sec. 3.3). The overview of our EvolveNav is presented in Figure 2. Specifically, our EvolveNav consists of two stages: 1) Formalized CoT Supervised Fine-Tuning (Sec. 3.3.1), where the LLM-based VLN model is initially trained using formalized CoT labels with supervised fine-tuning to activate the model’s navigational reasoning ability and simultaneously promote the reasoning speed; 2) Self-Reflective Post-Training (Sec. 3.3.2), where the model is further trained with its own reasoning outputs as self-enriched CoT labels to increase supervision diversity, accompanied by a self-reflective auxiliary task to encourage better learning of accurate navigational reasoning patterns by discriminating from incorrect ones.

### v3

# EvolveNav: Empowering LLM-Based Vision-Language Navigation via Self-Improving Embodied Reasoning

@@ Abstract
Recent studies have revealed the potential of training open-source Large Language Models (LLMs) to unleash LLMs’ reasoning ability for enhancing vision-language navigation (VLN) performance, and simultaneously mitigate the domain gap between LLMs’ training corpus and the VLN task. However, these approaches predominantly adopt straightforward input-output mapping paradigms, causing the mapping learning difficult and the navigational decisions unexplainable. Chain-of-Thought (CoT) training is a promising way to improve both navigational decision accuracy and interpretability, while the complexity of the navigation task makes the perfect CoT labels unavailable and may lead to overfitting through pure CoT supervised fine-tuning. To address these issues, we propose EvolveNav, a novel sElf-improving embodied reasoning paradigm that realizes adaptable and generalizable navigational reasoning for boosting LLM-based vision-language Navigation. Specifically, EvolveNav involves a two-stage training process: (1) Formalized CoT Supervised Fine-Tuning, where we train the model with curated formalized CoT labels to first activate the model’s navigational reasoning capabilities, and simultaneously increase the reasoning speed; (2) Self-Reflective Post-Training, where the model is iteratively trained with its own reasoning outputs as self-enriched CoT labels to enhance the supervision diversity. A self-reflective auxiliary task is also designed to encourage the model to learn correct reasoning patterns by contrasting with wrong ones. Experimental results under both task-specific and cross-task training paradigms demonstrate the consistent superiority of EvolveNav over previous LLM-based VLN approaches on various popular benchmarks, including R2R, REVERIE, CVDN, and SOON. EvolveNav open avenues for exploring effective self-improving reasoning paradigms, enabling building agents capable of self-evolving for promoting LLM-based embodied AI research.
publicationid: pubid: 0000–0000/00\$00.00 © 2021 IEEE
@@ 1 Introduction
Vision-Language Navigation (VLN) has received significant research interest within the Embodied AI community, due to its practicality and flexibility in enabling human-robot interaction in real-world robotic applications. In VLN tasks, an embodied agent needs to follow natural language instructions to navigate through complex visual environments to reach the target position. Early works improve VLN performance by designing dedicated model architectures [1, 2, 3, 4], introducing powerful learning paradigms [5, 6, 7], and developing useful data augmentation techniques [8, 9, 10, 11]. Subsequently, pretraining-based VLN approaches have been widely proposed to improve the cross-modal alignment ability and decision accuracy of navigation agents [12, 13, 14, 15, 16]. Nevertheless, constrained by the limited scale of pretraining and VLN in-domain data, these approaches cannot learn navigational reasoning knowledge sufficiently, and therefore still struggle to handle various unseen navigation scenarios.
With the rapid progress of large language models (LLMs) [17, 18, 19], emerging works have introduced LLMs to address embodied tasks by resorting to LLMs’ rich real-world common sense and powerful reasoning ability [20, 21, 22]. Some recent works have attempted to build LLM-based VLN models in a zero-shot or trainable manner  [23, 24, 25, 26, 27]. The zero-shot approaches, such as NavGPT [24] and MapGPT [23], resort to closed-source LLMs [28] to generate navigational reasoning and decision for different navigation timesteps. To alleviate the high cost of frequently querying closed-source LLMs for sequential decision making, the trainable methods collect in-domain data to train open-source LLMs [29] to build the navigation agent. However, they typically map directly from navigational inputs to decisions without explicit intermediate reasoning steps, leading to decision uninterpretability and may also limit performance (see Figure 1(a)).
@@
Recent studies have revealed the effectiveness of chain-of-thoughts (CoT) training in enhancing both the decision accuracy and interpretability for embodied tasks [30, 31, 32]. Most of these approaches employ the supervised fine-tuning (SFT) paradigm for conducting embodied CoT training. However, introducing CoT supervised fine-tuning for training VLN models is highly challenging due to the following two reasons. Firstly, due to the complexity and uncertainty of the navigation task, there can be multiple cues for deciding the correct navigation action, i.e., there may be no single “correct” CoT label to guide navigation for a specific timestep. This leads to a hard collection process of perfect navigation CoT supervision. Secondly, pure CoT supervised fine-tuning using fixed CoT labels may cause overfitting to certain reasoning patterns and thus harm the generalization to diverse unseen scenarios.
@@
In this paper, we propose a novel sElf-improving embodied reasoning paradigm for enhancing LLM-based vision-language Navigation, called EvolveNav, to fulfill generalizable and adaptable navigational reasoning under various tasks and scenarios. EvolveNav comprises two training phases: 1) Formalized CoT Supervised Fine-Tuning and 2) Self-Reflective Post-Training. The Stage 1 training of Formalized CoT Supervised Fine-Tuning aims to first activate the model’s potential reasoning capabilities, where we ask the model to produce explicit chain-of-thought navigational reasoning dynamically by predicting the landmarks needed to locate with the corresponding direction for deciding the navigation actions. To alleviate generating redundant reasoning and increase the inference speed, we conduct the CoT supervised fine-tuning using curated formalized CoT labels, which are collected by filling the landmark and direction information into concise label templates. Then, we conduct Self-Reflective Post-Training for Stage 2 training, aiming to mitigate the overfitting to pre-constructed CoT labels and enable self-improving reasoning for enhancing generalization. Specifically, we design a self-enriched CoT label enhancement scheme, where we train the model with its iteratively produced correct reasoning outputs to diversify the CoT supervision. We also construct a self-reflective auxiliary task, where the model needs to discriminate between positive and negative navigational reasoning to learn correct reasoning patterns. As shown in Figure 1, in contrast to CoT supervised fine-tuning using fixed CoT labels (i.e., Figure 1(b) Formalized CoT Learning and (c) Free-form CoT Learning), our EvolveNav (Figure 1(d)) can generate embodied CoT in a self-refining manner during training to mitigate the overfitting. Additionally, through training with formalized CoT labels, our EvolveNav can significantly reduce uninformative navigational reasoning to promote the reasoning speed compared with using free-form CoT labels for training (as in Figure 1 (c)).
@@
We conduct substantial experiments under both task-specific and cross-task training paradigms on multiple public VLN benchmarks, including R2R [33], CVDN [34], REVERIE [35], and SOON [36]. Experimental results show that EvolveNav significantly outperforms previous LLM-based VLN approaches on various benchmarks, demonstrating the effectiveness of our self-improving embodied reasoning paradigm in promoting navigation decision accuracy and generalization. We carefully conduct ablation experiments to explore how to design streamlined CoTs that can provide interpretability while boosting navigation performance. Visualization also insightfully reveals the reasonability of our design for CoT labels in enhancing decision interpretability and improving navigational reasoning.
@@
  • We propose EvolveNav, a novel self-improving embodied reasoning paradigm for enhancing LLM-based vision-and-language navigation, which fulfills generalizable and adaptable navigational reasoning under various tasks and scenarios.
@@
  • Experimental results demonstrate the superiority of EvolveNav over previous LLM-based approaches on various VLN benchmarks. Our EvolveNav can improve the generalization of both navigational reasoning and decision-making, providing meaningful insights for designing advanced embodied reasoning paradigms.
@@ 2.1 Vision-Language Navigation (VLN)
Vision-Language Navigation (VLN) has attracted intensive research interest in recent years. Various VLN benchmarks have been proposed to evaluate agents’ ability for navigational reasoning and instruction following [33, 37, 38, 34, 35, 36]. Previous approaches employ non-pretraining-based [1, 2, 3, 5, 8, 9] or pretraining-based paradigms [12, 13, 15, 16, 39, 40, 41] for tackling the above VLN tasks. However, these approaches cannot generalize well to diverse unseen scenarios that require rich real-world commonsense, and the navigation decisions also lack explainability. Some recent works have introduced LLMs to assist the VLN task, by either eliciting the useful navigation knowledge stored in LLMs [42, 43, 44] or employing the LLM as the navigation backbone for action decision [24, 45, 27, 32, 23]. Our work lies in the latter.
Different from previous approaches, we propose a new VLN framework in this work, where the LLM-based navigation backbone iteratively generates intermediate reasoning steps in a self-improving manner during training to guide navigational decisions. As a result, both the reasoning ability and decision interpretability of the navigation model can be significantly enhanced.
2.2 LLMs as Embodied Agents
@@
Recent research have revealed the giant potential of utilizing Large Language Models (LLMs) as embodied agents to complete the robotic navigation and manipulation tasks, benefiting from the outstanding ability of planning, reasoning, and reflection of LLMs [46, 20, 21, 47, 48, 49, 50, 51, 52]. For example, LM-Nav [48] introduces the LLM to parse the long navigation instruction into sequential landmarks for facilitating the navigational planning. Voxposer [51] introduces LLMs for code writing and combines them with the Vision-Language models (VLMs) to compose 3D value maps for robotic manipulation.
@@
There are typically two branches of works where the LLMs act as embodied agents for tackling the VLN task. In the first branch, closed-source LLMs like GPT-4 [53] are queried in a zero-shot manner to decide the action sequentially [24, 25, 23, 54]. For example, NavGPT [24] transforms visual observations into textual formats and feed them to LLMs for generating action predictions. The second branch finetunes open-source LLMs with in-domain VLN datasets, which alleviates the LLM’s query cost as well as mitigates the gap between LLM’s training corpus and VLN tasks [45, 32, 27, 55]. For example, Navid [27] constructs a video-based navigational vision-language model and train it using navigation samples collected from continuous R2R datasets. However, most of them map navigation inputs to action decisions directly without the reasoning output. In contrast, we train the open-source LLM to generate self-improving embodied reasoning explicitly to improve action decision accuracy, which enhances the decision interpretability as well as mitigates overfitting to training reasoning labels.
2.3 Embodied Chain-of-thoughts Training
@@
Chain-of-thoughts (CoT) reasoning [56] has been a widely utilized technique in Large Language Models (LLMs) and Vision-Language Models (VLMs). By generating the intermediate reasoning steps rather than directly predicting the answer, CoT can promote the answer accuracy for various tasks such as mathematical reasoning, commonsense reasoning, code generation, etc [53, 28, 57]. Inspired by this, some recent works have trained LLMs/VLMs to generate reasonable CoT for improving the action decision accuracy in embodied tasks [30, 32, 58, 31, 59, 43]. ECoT [30] trains vision-language-action models to generate embodied reasoning including object bounding boxes and end effector positions to encourage better adaptation for robotic manipulation tasks. NavGPT-2 [43] resorts to the GPT-4V model to collect free-form step-wise CoT reasoning data to improve the navigational reasoning ability of the VLN-specialized models [15]. CoT-VLA [31] incorporates explicit visual chain-of-thought reasoning into vision-language-action models by predicting future image frames before generating the action sequence.
In this work, we propose a novel self-improving embodied CoT training paradigm for boosting VLN performance, which effectively improves the reasoning ability and decision accuracy of the navigation model in various unseen scenarios. Moreover, we construct CoT supervision in a formalized manner, which significantly reduces redundant reasoning information and simultaneously promotes the model’s inference speed.
@@
Figure 2: Overview of EvolveNav. EvolveNav involves a two-phase training framework for fulfilling self-improving embodied reasoning. In Stage 1 Formalized CoT Supervised Fine-Tuning, the navigation agent is trained using pre-constructed formalized CoT labels to generate navigational reasoning by predicting the landmark needed to locate with the corresponding direction. In Stage 2 Self-Reflective Post-Training, the agent’s own reasoning outputs are introduced as the self-enriched CoT labels to enhance the supervision diversity. A self-reflective auxiliary task is also designed to guide the navigation agent to discriminate between correct and wrong reasoning outputs.
@@ 3 Method
In this section, we first introduce the problem definition of the VLN task (Sec. 3.1). Then, we present the model architecture of the LLM-based navigation agent in our EvolveNav (Sec. 3.2). Finally, we delve into the details of the proposed self-improving embodied reasoning framework (Sec. 3.3).
The overview of our EvolveNav is presented in Figure 2. Specifically, EvolveNav consists of two training stages: 1) Formalized CoT Supervised Fine-Tuning (Sec. 3.3.1), where we curate formalized CoT labels to initially train the LLM-based VLN model with supervised fine-tuning to activate the model’s navigational reasoning ability and simultaneously promote the reasoning speed; 2) Self-Reflective Post-Training (Sec. 3.3.2), where the model is further trained with its own reasoning outputs as self-enriched CoT labels to increase supervision diversity, accompanied by a self-reflective auxiliary task to encourage better learning of accurate navigational reasoning patterns by discriminating from incorrect ones.